OpenStudy (bronzbattalion):

1. A recipe calls for cups of flour, cup of white sugar, and cup of brown sugar. The recipe makes 6 servings. (a) How many cups of flour are there per serving? Show your work. (b) How many total cups of sugar (white and brown) are there per serving? Show your work. (c) Suppose you modify the recipe so that it makes 9 servings. How much more flour do you need for the modified recipe than you need for the original recipe? Show your work.

OpenStudy (bronzbattalion):

@KJSaif:

OpenStudy (bronzbattalion):

@johnp12

OpenStudy (bronzbattalion):

@just_one_last_goodbye

OpenStudy (acannell):

Flour = 1/1/2 white = 3/4 brown = 1/3 servings = 6 a) 6 ÷ 1/1/2 = 4 cups b) w = 6 ÷ 3/4 w = 8 cups b = 6 ÷ 1/3 b = 18 cups total = 8 + 18 total = 26 cups c) 9 ÷ 1/1/2 = 6 cups 6 - 4 = 2 cups more Now i know that this is correct!!! :)